Aaleighacruz:

A cone has a diameter of 4 centimeters and a height that is 8 times the diameter. Using 3.14 for pi, which of the following can be used to calculate the volume of the cone? (6 points)

josiahcantcount:

I dont see the options

Aaleighacruz:

one over three (3.14)(32cm)2(4cm) one over three (3.14)(32cm)2(8cm) one over three (3.14)(4cm)2(32cm) one over three (3.14)(2cm)2(32cm)

jhonyy9:

do you know formula for volume of a cone ?

Aaleighacruz:

no

josiahcantcount:

The formula for the volume of a cone is V=1/3hπr².

jhonyy9:

V = area of base time height divide 3

jhonyy9:

area of base = ?

Aaleighacruz:

idk im sorry

jhonyy9:

area of base = area of a circle the radius of this circle = diameter /2 diameter = 4 cm

Aaleighacruz:

would it be A?

AZ:

No Look at the information they gave you Diameter = 4 Height is 8 times the diameter so 8*4 = ?? You need to know the formula for the volume of a cone \(\text{Volume} = \dfrac{1}{3} \pi r^2 h\) \(\pi\) or pi is 3.14 r is the radius of the base of the cone Remember that radius = 1/2 * diameter The radius is half of the diameter. So if the diameter is 4, what is the radius?? And h is height Put all the numbers you get into the formula

Aaleighacruz:

so its C?

AZ:

If the diameter is 4 Then what is the radius? radius = 1/2 * diameter

AZ:

What is half of 4?

Aaleighacruz:

2

AZ:

So your radius is 2 The formula for volume has RADIUS not diameter So can you replace r with 2 What's your final answer?

Aaleighacruz:

1/3 (3.14)(2cm)2(32cm)

AZ:

That is correct

Aaleighacruz:

thank you so much

AZ:

You're welcome!